### v1

# A constant rank theorem for linear elliptic equations on the sphere with applications to the mixed Christoffel problem

Abstract.
We study the mixed Christoffel problem for $C^{2,+}$ convex bodies providing sufficient conditions for its solution. Key to our approach is a constant rank theorem, following the approach developed in [13] to address the Christoffel problem, in order to ensure that the solution to a related second order linear PDE on the sphere is indeed geometric, that is, it is the support functions of a $C^{2,+}$ convex body.
Key words and phrases:
convex body; mixed are measures; mixed Christoffel problem; elliptic PDE; constant rank theorem
2000 Mathematics Subject Classification:
52A20, 35J15
1. Introduction
In this paper we are interested in (mixed) area measures of convex bodies (compact and convex subsets of $\mathbb{R}^{n+1}$, with non-empty interior). Let us briefly explain how these measures are generated. It is well known that the volume (that is, the Lebesgue measure, here denoted by $V$) of the Minkowski linear combination
$t_{1}\Omega_{1}+\dots+t_{m}\Omega_{m},\quad t_{1},\dots,t_{m}\geq 0,$
of $m$ convex bodies in $\mathbb{R}^{n+1}$ is an homogeneous polynomial of degree $(n+1)$ in the variables $t_{i}$’s:
$V(t_{1}\Omega_{1}+\dots+t_{m}\Omega_{m})=\sum_{i_{1},\dots,i_{n+1}=1}^{m}V(\Omega_{i_{1}},\dots,\Omega_{i_{n+1}})t_{i_{1}}\cdots t_{i_{n+1}}.$
The coefficients of this polynomial, $V(\Omega_{i_{1}},\dots,\Omega_{i_{n+1}})$, are the so-called mixed volumes. Hence, each mixed volume is a quantity depending on $(n+1)$ convex bodies:
$V=V(\Omega_{1},\dots,\Omega_{n+1}).$
Mixed volumes are among the fundamental notions in convex geometry (see for instance [21]). They include quermassintegrals as special cases: for a fixed convex body $\Omega$, mixed volumes of the form
$V(\underbrace{\Omega,\dots,\Omega}_{\mbox{\tiny$k$-times}},\underbrace{B_{1},\dots,B_{1}}_{\mbox{\tiny$(n+1-k)$-times}})$
where $B_{1}$ is the unit ball of $\mathbb{R}^{n+1}$ and $k\in\{1,\dots,n+1\}$, are called quermassintegrals of $\Omega$ (or, with a different normalization, its intrinsic volumes).
Mixed area measures arise naturally in integral representations of mixed volumes. Given $n$ convex bodies $\Omega_{1},\dots,\Omega_{n}$ in $\mathbb{R}^{n+1}$, there exists a finite Borel measure $S=S(\Omega_{1},\dots,\Omega_{n};\cdot)$, defined on the unit sphere ${\mathbb{S}}^{n}$} of $\mathbb{R}^{n+1}$, such that, for any other convex body $\Omega$,
$V(\Omega_{1},\dots,\Omega_{n},\Omega)=\int_{{{\mathbb{S}}^{n}}}u_{\Omega}(x)dS(\Omega_{1},\dots,\Omega_{n};x),$
where $u$ is the support function of $\Omega$. We recall that the support function of $\Omega$ (with respect to the origin) is the function $u_{\Omega}:\ {{\mathbb{S}}^{n}}\to\mathbb{R}$ defined by
$u_{\Omega}(x)=\sup_{z\in\Omega}\langle x,z\rangle,$
where $\langle\cdot,\cdot\rangle$ is the standard scalar product in $\mathbb{R}^{n+1}$. The measure $S=S(\Omega_{1},\dots,\Omega_{n};\cdot)$ is called the mixed area measure of $\Omega_{1},\dots,\Omega_{n}$. Due to translation invariance of mixed volumes, area measures verify the following condition:
(1) $\int_{\mathbb{S}^{n}}x_{j}dS(\Omega_{1},\dots,\Omega_{n};x)=0,\quad\forall j=1,\cdots,n+1$
(here $x_{1},\dots,x_{+1}$ are the coordinates in $\mathbb{R}^{n+1}$).
A special class of mixed area measures related to a convex body $\Omega$ are its area measures, $S_{k}=S_{k}(\Omega;\cdot)$, $k=1,\dots,n$, defined by
$S_{k}(\Omega;\cdot)=S(\underbrace{\Omega,\dots,\Omega}_{\mbox{\tiny$k$-times}},\underbrace{B_{1},\dots,B_{1}}_{\mbox{\tiny$(n-k)$-times}};\cdot).$
In particular, the measure $S_{n}(\Omega;\cdot)$ has the following simple geometric interpretation: for every Borel set $\eta\subset{{\mathbb{S}}^{n}}$,
$S_{n}(\Omega;\eta)={\mathcal{H}}^{n-1}(\nu_{\Omega}^{-1}(\eta)),$
where $\mathcal{H}^{n-1}$ is the (n1)(n-1)-dimensional Hausdorff measure, and $\nu_{\Omega}$ is the Gauss map of $\Omega$.
The celebrated Minkowski problem requires to find a convex body $\Omega$ with prescribed area measure $S_{n}(\Omega;\cdot)$. This problem is well-posed: simple necessary and sufficient conditions on the given measure, in order to have existence and uniqueness (up to translations) of the solution, are known, and a satisfactory regularity theory has now been developed, thanks to the contributions of many authors. The literature concerning this problem is vast; the interested reader is referred, for instance, to Chapter 7 in [21], for an exhaustive account on this subject.
On the other hand, the problem of finding a convex body with given area measure of order one, that is $S_{1}(\Omega;\cdot)$, called the Christoffel problem, appears to be much more delicate. Necessary and sufficient conditions for the existence of a solution were established by Firey [12] and Berg [3] (see also Chapter 8 in [21] and [19]), but they are in general not easy to be checked in practice. The situation is even more complicated for the Christoffel-Minkowski problem, concerning the area measures of intermediate order, $S_{k}(\Omega;\cdot)$, $2\leq k\leq(n-1)$.
1.1. The passage to PDE’s
A general observation, which is crucial for the present paper, is that, under regularity assumptions on the involved convex bodies, mixed area measures admit a density with respect to the uniform measure on ${\mathbb{S}}^{n}$}, and such densities can be explicitly written in terms of the derivatives of the support function. This allows, for instance, to reduce the Minkowski and the Christoffel problems to find a solution of a partial differential equation on the unit sphere, within the class of support functions of convex bodies.
We say that a convex body $\Omega$ is of class $C^{2,+}$ if its boundary is of class $C^{2}$, and the Gauss curvature is positive at every point of $\partial\Omega$. If $\Omega$ is of class $C^{2,+}$, then its support function $u$ belongs to $C^{2}({{\mathbb{S}}^{n}})$. We denote by $u_{i}$, $u_{ij}$, $i,j=1,\dots,n$, the first and second covariant derivatives of $u$, with respect to a local orthonormal frame on ${\mathbb{S}}^{n}$}.
Definition 1.1.
For a convex body $\Omega\subset\mathbb{R}^{n+1}$, of class $C^{2,+}$, with support function $u$, we define the inverse Weingarten form of $\Omega$ as
(2) $W_{\Omega}(x)=(u_{ij}(x)+u(x)\delta_{ij})_{i,j=1,\dots,n},\quad x\in{{\mathbb{S}}^{n}}.$
Note that $W$ is independent of the choice of the point with respect to which the support function is defined (that is, is invariant with respect to translations of $\Omega$). Moreover, $\Omega$ is of class $C^{2,+}$ if and only if its inverse Weingarten form is positive definite at every point of ${\mathbb{S}}^{n}$}; see, for instance, Chapter 2 in [21].
We will denote by $\mathscr{D}$ the mixed discriminant operator, that is, the operator generated by the polarization of the determinant with respect to the usual addition of matrices. $\mathscr{D}$ acts on $n$-tuples of $n\times n$ symmetric matrices. With this notation, if $\Omega_{1},\dots,\Omega_{n}$ are convex bodies of class $C^{2,+}$, and $W_{i}$ is the inverse Weingarten form of $\Omega_{i}$, $i=1,\dots,n$, then their mixed area measure is absolutely continuous with respect to the (n1)(n-1)-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}, and its density is
(3) $dS(\Omega_{1},\dots,\Omega_{n};x)=\mathscr{D}(W_{1},\cdots,W_{n})dx.$
In particular
$dS_{n}(\Omega;x)=\det(W_{\Omega}(x))dx$
for every convex body $\Omega$ of class $C^{2,+}$.
We now observe that the support function of the unit ball $B_{1}$ of $\mathbb{R}^{n+1}$ (restricted to ${\mathbb{S}}^{n}$}) is the constant function $1$, and consequently
$W_{B_{1}}\equiv I_{n},$
where $I_{n}$ is the identity matrix of order $n$. Hence
$dS_{k}(\Omega;x)=\mathscr{D}(\underbrace{W_{\Omega}(x),\dots,W_{\Omega}(x)}_{\mbox{\tiny$k$-times}},\underbrace{I_{n},\dots,I_{n}}_{\mbox{\tiny$(n-k)$-times}})dx=\sigma_{k}(W_{\Omega}(x))dx,$
where $\sigma_{k}(W_{\Omega})$ denotes the $k$-th principal symmetric function of the eigenvalues of $W_{\Omega}$.
As a consequence, the Minkowski problem in the smooth setting can be reduced to solve a partial differential equation on ${\mathbb{S}}^{n}$}: if $f$ is the density of the given measure, the problem amounts to find a solution $u$ to the Monge-Ampère equation
$\det(u_{ij}(x)+u\delta_{ij}(x))=f(x),$
such that its associated inverse Weingarten form is positive definite on ${\mathbb{S}}^{n}$}:
(4) $W(x)=(u_{ij}(x)+u\delta_{ij}(x))>0\quad\forall\,x\in{{\mathbb{S}}^{n}}.$
The latter condition ensures that $u$ is the support function of a convex body (of class $C^{2,+}$). We will refer to this condition saying that $u$ is a geometric solution of the relevant equation.
Similarly, the Christoffel problem is equivalent to find a geometric solution of the linear equation
(5) $\mbox{tr}(W)=\Delta_{\mathbb{S}^{n}}u+nu=f,$
where $\Delta_{{{\mathbb{S}}^{n}}}$ denotes the spherical Laplacian. More generally, for an arbitrary $k\in\{1,\dots,n\}$, the Christoffel-Minkowski problem is reduced to find a geometric solution of the following Hessian equation
(6) $\sigma_{k}(u_{ij}+u\delta_{ij})=f.$
Here, given a symmetric matrix $W$, $\sigma_{k}(W)$ denotes the $k$-th elementary symmetric function of the eigenvalues of $W$. In all cases, in view of (1), the compatibility condition
(7) $\int_{\mathbb{S}^{n}}x_{j}f(x)dx=0,\quad\forall\ j=1,\dots,n+1,$
is a necessary condition on $f$. This condition is verified, for instance, when $f$ (or equivalently $\mu$) is even. In the case of Minkowski problem (in the smooth setting), (7) is also sufficient. On the other hand it is known that such condition is not sufficient for the existence of a solution of the Christoffel problem (see [21, Section 8.3]).
1.2. A sufficient condition for the Christoffel problem
Existence of solution to (5) follows from the standard theory of elliptic PDE’s. The main question is when such solution $u$ is geometric, i.e. $u$ is a support function of convex body. A progress was made in the paper [13], where the authors established the following constant rank type theorem, for solutions of equation (5), under a fairly simple condition on $f$ (in the same paper they proved a similar result for equation (6)).
Theorem A.
Let $f\in C^{1,1}({{\mathbb{S}}^{n}})$ be a strictly positive function satisfying (7). Assume moreover that
(8) the $1$-homogeneous extension of $\dfrac{1}{f}$ to $\mathbb{R}^{n+1}$, is convex.
If $u$ is a solution to (5) with
$W(x)=(u_{ij}(x)+u(x)\delta_{ij})\geq 0\quad\forall\,x\in{{\mathbb{S}}^{n}},$
then $W(x)>0$ for every $x\in\mathbb{S}^{n}$.
Using the previous result and standard continuity argument, existence of a geometric solution to (5) can be proved if there is a homotopic path, described by a real parameter $t\in[0,1]$, connecting $f$ and, say, the constant function $1$, such that the assumption (8) is valid for every $t$. For that purpose, Andrews and Ma used a curvature flow in an unpublished note [2]; see also a recent work by Bryan-Ivaki-Scheuer [7]. A direct homotopic path was constructed by Sheng-Trudinger-Wang in [22]. Therefore, as a corollary of Theorem A, conditions (1) (that is (7)) and (8) are sufficient for the existence of a solution of the Christoffel problem. Note that (1) is also a necessary condition.
1.3. The mixed Christoffel problem
In this paper we study the following problem.
Problem 1.2 (The mixed Christoffel problem).
Let $\Omega_{1},\dots,\Omega_{n-1}$, be given convex bodies in $\mathbb{R}^{n+1}$, and let $\mu$ be a Borel measure on ${\mathbb{S}}^{n}$}. Find a convex body $\Omega$ in $\mathbb{R}^{n+1}$, such that
(9) $S(\Omega_{1},\dots\Omega_{n-1},\Omega;\cdot)=\mu(\cdot).$
This problem is mentioned in the recent paper [15]. Moreover, necessary and sufficient conditions for the existence of a solution in the rotation invariant case have been recently established in [5, 6, 20].
If $\Omega_{i}=B_{1}$ for every $i=1,\dots,(n-1)$, then the problem reduces to the Christoffel problem. When $L$ is a given convex $C^{2,+}$ set (different from $B_{1}$) and $\Omega_{i}=L$ for every $i=1,\dots,(n-1)$, then the problem can be interpreted as an anisotropic Christoffel problem, where $L$ plays the role of the unit ball in place of $B_{1}$.
We first note that, due to (1), the following is a necessary condition for the existence of a solution:
(10) $\int_{\mathbb{S}^{n}}x_{j}d\mu(x)=0,\quad\forall j=1,\cdots,n+1.$
We will treat the mixed Christoffel problem in the smooth setting: we will assume that $\Omega_{1},\cdots,\Omega_{n-1}$ are convex bodies of class $C^{2,+}$, and that the measure $\mu$ has a smooth density $f$ with respect to the $n$-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}. In particular, condition (10) is equivalently expressed by (7).
We establish a sufficient condition involving $\Omega_{1},\dots,\Omega_{n-1}$ and $f$, for the existence of a solution to the mixed Christoffel problem.
In order to state our existence result, we need to recall the notion of mixed co-factor matrix. Let $M_{1},\dots,M_{n-1}$ be $n\times n$ symmetric matrices. For any further $n\times n$ symmetric matrix $M=(m_{ij})_{i,j=1\dots,n}$, the mixed discriminant
$\mathscr{D}(M_{1},\dots,M_{n-1},M)$
is linear with respect to $M$. Therefore it can be written in the form
$\mathscr{D}(M_{1},\dots,M_{n-1},M)=\sum_{i,j=1}^{n}c_{ij}m_{ij},$
for suitable and uniquely determined coefficients $c_{ij}\in\mathbb{R}$. The matrix
$C(M_{1},\dots,M_{n-1})=(c_{ij})_{i,j=1,\dots,n},$
is called the mixed co-factor matrix of $M_{1},\dots,M_{n-1}$. Alternatively, the coefficients $c_{ij}$ can be defined as
(11) $c_{ij}=\frac{\partial\mathscr{D}(M_{1},\dots,M_{n-1},M)}{\partial m_{ij}},\quad\forall\,i,j=1,\dots,n.$
By (3), condition (9) is equivalent to
(12) $\mathscr{D}(W_{1},\cdots,W_{n-1},W)=f,\quad\mbox{on $\mathbb{S}^{n}$},$
and the latter equation can be written in the form:
(13) $\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}(x)(u_{\alpha\beta}(x)+\delta_{\alpha\beta}u(x))=f(x)$
(once that a local orthonormal coordinate system has been fixed on ${\mathbb{S}}^{n}$}), where
$(a_{\alpha\beta})_{\alpha,\beta=1,\dots,n}$
is the mixed co-factor matrix of $W_{1},\cdots,W_{n-1}$. Equation (13) is a linear second order equation in the unknown function $u$. The fact that $\Omega_{i}$ is of class $C^{2,+}$ for every $i$, implies that this equation is uniformly elliptic, which is equivalent to
(14) $(a_{\alpha\beta}(x))>0,\quad\forall\ x\in{{\mathbb{S}}^{n}}.$
The previous conditions guarantee the existence of a solution to (13). Moreover, the solution is unique modulus the addition of linear functions (see, for instance, [14]). We also observe that if $\partial\Omega_{1},\cdots,\partial\Omega_{n-1}\in C^{4}$, then $a_{\alpha\beta}\in C^{2}$. Under these assumptions, the admissible solution $u$ of (13) belongs to $C^{4,\gamma}$ (but note that $C^{3,1}$ would be sufficient for us to carry out convexity estimates in later sections).
We look for a solution $u$ of (13) such that condition (4) holds. As in [13], we will prove a constant rank theorem for (13). In order to state our result, we need the following remark.
Remark 1.3.
Assume that $A$ is a symmetric $(2,0)$ tensor defined on an open subset $D$ of ${\mathbb{S}}^{n}$}, and that $u\in C^{2}(D)$. For a fixed local coordinate system on ${\mathbb{S}}^{n}$}, let
$A=(a_{ij})_{i,j=1,\dots,n},\quad W=(w_{ij})_{i,j=1,\dots,n}=(u_{ij}+u\delta_{ij})_{i,j=1,\dots,n}.$
Then the quantity
${\rm tr}(AW)=\sum_{i,j=1}^{n}a_{ij}w_{ij},$
is independent of the local coordinate system chosen on ${\mathbb{S}}^{n}$}.
In the following statement $(a^{\alpha\beta})_{\alpha,\beta=1,\dots,n}$ denotes the inverse matrix of $(a_{\alpha\beta})_{\alpha,\beta=1,\dots,n}$.
Theorem 1.1.
Let $A(x)=(a_{ij}(x))\in C^{2,\gamma}(\mathbb{S}^{n})$ be a positive $(2,0)$-tensor in an open and connected set $D\subset\mathbb{S}^{n}$ and let $f\in C^{2,\gamma}(D)$ for some $\gamma>0$, with $f>0$ on $\mathbb{S}^{n}$. Suppose that, for every $1\leq l\leq n-1$, and every $q=1,\dots,n$,
(15) $\left(\delta_{ij}+\frac{\nabla_{i}\nabla_{j}(\frac{a_{qq}}{f})}{\frac{a_{qq}}{f}}+\frac{1}{2}\frac{\nabla_{i}a_{qq}\nabla_{j}a_{qq}}{a^{2}_{qq}}-\frac{1}{2}\sum_{\alpha,\beta=1}^{l}\frac{a^{\alpha\beta}\nabla_{i}a_{q\alpha}\nabla_{j}a_{q\beta}}{a_{qq}}\right)_{i,j=1,\dots,n}\geq 0.$
Let $u\in C^{2}(D)$ be a solution of equation (13) in $D$. If
$W=(u_{ij}+u\delta_{ij})_{i,j=1,\dots,n}\geq 0$
in $D$, then $W$ has constant rank in $D$. Moreover, if $D=\mathbb{S}^{n}$ (and $n\geq 2$), then $W$ has maximal rank, that is $W>0$ in $D$.
In the case $n=2$, condition (15) can be written as follows: for every $q=1,2$,
(16) $\left(\frac{a_{qq}}{f}\delta_{ij}+\nabla_{i}\nabla_{j}\left(\frac{a_{qq}}{f}\right)\right)_{i,j=1,2}\geq 0.$
Thank to (16), for $n=2$ a simplified version of the previous result can be achieved. We need to introduce a definition, first. For an arbitrary $k\in\mathbb{N}$, we denote by $\mbox{\rm Sym}(k)$ the space of square and symmetric matrices of order $k$.
Definition 1.4.
Let $n,k\in\mathbb{N}$, and let
$M\colon\mathbb{R}^{n+1}\to\mbox{\rm Sym}(k).$
We say that the function $M$ is convex if
$M((1-t)x_{0}+tx_{1})\leq(1-t)M(x_{0})+tM(x_{1}),$
for every $x_{0},x_{1}\in\mathbb{R}^{n+1}$ and for every $t\in[0,1]$. The above inequality has to be intended in the matricial sense: the difference between the right hand side and the left hand side is a positive semi-definite matrix.
Theorem 1.2.
Let $A(x)=(a_{ij}(x))_{i,j=1,2}\in C^{2,\gamma}(D)$ be a positive $(2,0)$-tensor defined in an open connected set $D\subset\mathbb{S}^{2}$, and let $f\in C^{2,\gamma}(D)$ for some $\gamma>0$, with $f>0$ on $D$. Let $M$ be the $1$-homogeneous extension of the tensor
$\frac{A}{f}$
to the cone
$C=\left\{x\in\mathbb{R}^{3}\colon\frac{x}{|x|}\in D\right\}.$
Assume that
(17) $M\mbox{ is locally convex in $C$}.$
Let $u\in C^{2}(\mathbb{S}^{2})$ be a solution of equation (13). If
$W=(u_{ij}+u\delta_{ij})_{i,j=1,2}\geq 0$
in $D$, then $W$ has constant rank in $D$. In particular, if $D=\mathbb{S}^{2}$ then $W$ has full rank in $\mathbb{S}^{2}$.
Using the previous results, Theorem A and a continuity argument, we will deduce the following full rank theorem.
Corollary 1.5.
Under the same assumptions of Theorem 1.1 (resp. Theorem 1.2) on AA and $f$, with $D={{\mathbb{S}}^{n}}$, assume moreover that $f$ verifies (8). Let $u$ be a solution of equation (13). Then
$W=(u_{ij}+u\delta_{ij})>0\quad\mbox{on ${{\mathbb{S}}^{n}}$.}$
As an application of Corollary 1.5, we get the following existence result for the mixed Minkowski problem.
Theorem 1.3.
Let $\Omega_{1},\dots,\Omega_{n}$ be convex bodies in $\mathbb{R}^{n+1}$, with boundary of class $C^{4,\gamma}$, for some $\gamma>0$, and positive Gauss curvature. For every $r=1,\dots,n$, let $u_{r}$ be the support function of $\Omega_{r}$ and
$W_{r}(x)=((u_{r})_{ij}(x)+u_{r}(x)\delta_{ij})_{i,j=1\dots,n},\quad x\in{{\mathbb{S}}^{n}},$
be the inverse Weingarten form of $\Omega_{r}$. Let
$A(x)=(a_{ij}(x))$
be the mixed co-factor matrix of $W_{1}(x),\dots,W_{n-1}(x)$. Let $f\in C^{2,\gamma}(\mathbb{S}^{n})$, with $f>0$ on $\mathbb{S}^{n}$, and assume that $f$ verifies condition (7) and condition (8). Let $\mu$ be the Borel measure on ${\mathbb{S}}^{n}$} with density $f$ with respect to the (n1)(n-1)-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}.
Suppose that, for any $1\leq l\leq n-1$, and any $q=1,\ldots,n$,
(18) $\left(\delta_{ij}+\frac{\nabla_{i}\nabla_{j}(\frac{a_{qq}}{f})}{\frac{a_{qq}}{f}}+\frac{1}{2}\frac{\nabla_{i}a_{qq}\nabla_{j}a_{qq}}{a^{2}_{qq}}-\frac{1}{2}\sum_{\alpha,\beta=1}^{l}\frac{a^{\alpha\beta}\nabla_{i}a_{q\alpha}\nabla_{j}a_{q\beta}}{a_{qq}}\right)_{i,j=1,\dots,n}\geq 0.$
Then there exists a convex body $\Omega$ in $\mathbb{R}^{n+1}$, of class $C^{2,+}$, such that
$S(\Omega_{1},\dots,\Omega_{n-1},\Omega;\cdot)=\mu(\cdot).$
Moreover, $\Omega$ is uniquely determined up to translations.
In the case $n=2$, condition (18) reads as (16).
Remark 1.6.
If $A(x)=(a_{ij}(x))$ is the identity matrix for every $x\in{{\mathbb{S}}^{n}}$, that is, we are in the case of the ordinary Christoffel problem, condition (18) becomes
$\left(\left(\frac{1}{f}\right)_{ij}+\delta_{ij}\frac{1}{f}\right)_{ij}\geq 0.$
In other words, we get the sufficient condition (8) found by Guan and Ma in [13].
Remark 1.7.
Condition (8) is obviously verified if $f\equiv 1$.
1.4. The mixed Christoffel problem in the $3$-dimensional space
The $3$-dimensional case ($n=2$), deserves a special treatment. The mixed Christoffel problem in $\mathbb{R}^{3}$ can be formulated as follows: let $\Omega_{0}$ be a convex body in $\mathbb{R}^{3}$, and let $\mu$ be a Borel measure on $\mathbb{S}^{2}$, such that
$\int_{\mathbb{S}^{2}}xd\mu(x)=0.$
The problem requires to find a convex body $\Omega$ such that
$S(\Omega_{0},\Omega;\cdot)=\mu(\cdot).$
Assume that $\Omega_{0}$ is of class $C^{2,+}$, with support function $h\in C^{2}(\mathbb{S}^{2})$; let
$W_{0}=W_{\Omega_{0}}=(h_{ij}+h\delta_{ij})_{i,j=1,2},$
be the inverse Weingarten map of $\Omega_{0}$. Assume moreover that $\mu$ has a density $f\in C^{2,\gamma}(\mathbb{S}^{2})$, for some $\gamma\in(0,1)$.
Using Theorem 1.2, or equivalently (16), we will derive two sufficient conditions for the existence of a solution, stated in the following Theorems 1.4 and 1.5.
Theorem 1.4.
Let $\Omega_{0}\subset\mathbb{R}^{3}$ be a convex body with boundary of class $C^{4,\gamma}$, for some $\gamma>0$, and with positive Gauss curvature. Let $h$ be the support function of $\Omega_{0}$. Let $f\in C^{2,\gamma}(\mathbb{S}^{2})$, with $f>0$ on $\mathbb{S}^{2}$, be such that condition (8) holds. Assume that, for every $x\in\mathbb{S}^{2}$, and for every tangential direction $\alpha$ to $\mathbb{S}^{2}$ at $x$, the following inequality holds:
(19) $\left(\frac{h_{\alpha\alpha}(x)+h(x)}{f(x)}\right)_{\alpha\alpha}+\frac{h_{\alpha\alpha}(x)+h(x)}{f(x)}\geq 0,$
where sub-indices denote covariant partial derivatives with respect to the direction $\alpha$. Then, there exists a convex body $\Omega\in C^{2,+}$ such that
$S(\Omega_{0},\Omega,\cdot)=\mu(\cdot)\,$,
where $\mu$ is the measure such that $d\mu(x)=f(x)dx$.
Example 1.8.
Let us show an example of application of the previous result. Let $f\equiv 1$ so that, in particular, condition (8) is verified. In this case (19) becomes:
(20) $h_{\alpha\alpha\alpha\alpha}+2h_{\alpha\alpha}+h>0,\quad\mbox{on $\mathbb{S}^{2}$.}$
Assume that $h\colon{{\mathbb{S}}^{n}}\to\mathbb{R}$ is of the form
$h=C+\psi$
where $C>0$ and $\psi\in C^{4}(\mathbb{S}^{2})$ verifies
$\|\psi\|_{C^{4}(\mathbb{S}^{2})}<\frac{C}{4}.$
Then (20) is verified.
Theorem 1.5.
Let $\Omega_{0}\subset\mathbb{R}^{3}$ and $f$ be as in Theorem 1.4, and let $W_{0}$ be the inverse Weingarten map of $\Omega_{0}$. If $f$ verifies condition (8), and the $1$-homogeneous extension of the tensor
$\frac{W_{0}}{f}$
is convex in $\mathbb{R}^{3}$, then there exists a convex body $\Omega\in C^{2,+}$ such that
$S(\Omega_{0},\Omega,\cdot)=\mu(\cdot),$
where $\mu$ is the Borel measure on $\mathbb{S}^{2}$ such that $d\mu=fdx$.
Remark 1.9.
In the assumptions of Theorem 1.5, let $h$ be the support function of $\Omega_{0}$, extended as a $1$-homogeneous function to $\mathbb{R}^{3}$. Let $D^{2}h(x)$ denote the (Euclidean) Hessian matrix of $h$, for every $x\neq 0$. $D^{2}h$ is a matrix-valued function defined in $\mathbb{R}^{n}\setminus{0}$, which extends the tensor $W_{0}$. Moreover, as $h$ is $1$-homogeneous, $D^{2}h$ is 1-1 homogeneous. As a consequence
$x\,\rightarrow\,|x|^{2}D^{2}h(x),$
is a $1$-homogeneous extension of $W_{0}$. Hence
$x\,\rightarrow\,|x|^{2}\left[f\left(\frac{x}{|x|}\right)\right]^{-1}D^{2}h(x),$
is the $1$-homogeneous extension of
$\frac{W_{0}}{f}.$
Therefore, the sufficient condition of Theorem 1.5, for the existence of a solution to the mixed Christoffel problem is equivalent to the convexity of the matrix-valued map $M\colon\mathbb{R}^{n}\to{\rm Sym}(n+1)$ defined by
$M(x)=|x|^{2}\left[f\left(\frac{x}{|x|}\right)\right]^{-1}D^{2}h(x).$
The rest of the paper is divided in two distinct and almost independent parts. Section 2.1 is devoted to the three-dimensional case ($n=2$), while in the final section we consider the case $n\geq 3$.
Acknowledgments. A. Colesanti was supported by the project Disuguaglianze analitiche e geometriche, funded by the Gruppo per l’Analisi Matematica la Probabilità e le loro Applicazioni. The research of the P. Guan is supported in part by an NSERC Discovery grant. M. Focardi has been supported by the European Union - Next Generation EU, Mission 4 Component 1 CUP B53D2300930006, codice 2022J4FYNJ, PRIN2022 project “Variational methods for stationary and evolution problems with singularities and interfaces”. P. Salani has been supported by the European Union - Next Generation EU, Mission 4 Component C2, PRIN2022 project “Geometric-Analytic Methods for PDEs and Applications (GAMPA)”, codice 2022SLTHCE, CUP B53D23009460006. M. Focardi and P, Salani are members of GNAMPA - INdAM.
The project was initiated in 2008 and at the time we were only be able to deal with $n=2$ case. The general dimensional case was worked out very recently. P. Guan author would like to thank the department of Mathematics at the University of Florence for the hospitality.
2. The case $n=2$
2.1. Proof of Theorem 1.1 in the case $n=2$
The proof is based on the constant rank method. We start with an observation which will be used also in the general case: using equation (13), the fact that the tensor $A$ is positive definite, and the condition $f>0$, we deduce that the rank of $W$ can not be 0. In particular, as $n=2$, the rank of $W$ can be $1$ or $2$.
If the rank equals $2$ at every point, there is nothing to prove.
Hence, to proceed we assume that there exists a point $x_{0}$ where the rank is $1$. In a sufficiently small neighborhood $\mathcal{U}$ of $x_{0}$, one of the eigenvalues of $W$ is larger than an absolute positive constant $C_{0}$, while the other one is close to zero, and can be assumed to belong to $0,C_{0}/2]$, that is,
$0\leq\lambda_{1}\leq\frac{C_{0}}{2}<C_{0}\leq\lambda_{2}\quad\text{in }\mathcal{U}\,$,
where, for $z\in\mathcal{U}$, $\lambda_{1},\lambda_{2}$ denotes the eigenvalues of $W(z)$. Let us notice that, by the standard regularity theory for elliptic equations, the solution $u\in C^{4,\gamma}(D)$ for some $\gamma\in(0,1)$. Then $\det(W),\lambda_{1}$ and $\lambda_{2}$ are of class $C^{2}({\mathcal{U}})$. We set
(21) $\phi(x)=\det(W(x))=w_{11}(x)w_{22}(x)-w_{12}^{2}(x),\quad x\in{\mathcal{U}}.$
We now fix $z\in{\mathcal{U}}$ and choose a local orthonormal frame of $\mathbb{S}^{2}$, such that $W(z)$ is diagonal: note that this is possible since, by our assumptions, the quantity
$\sum_{i,j=1}^{2}a_{ij}w_{ij}=\sum_{i,j=1}^{2}a_{ij}(u_{ij}+u\delta_{ij})$
is independent of the local coordinate system. In the next relations, all the involved functions will be computed at $z$. From the definition of $\phi$ we have
$w_{11}=\frac{1}{w_{22}}\phi,$
so that
(22) $0\leq w_{11}\leq\frac{1}{C_{0}}\phi.$
In the sequel we will denote by $C$ a general constant independent of $z\in\mathcal{U}$, and depending possibly only on $\|u\|_{C^{3}({\mathcal{U}})}$, $\|\phi\|_{C^{2}({\mathcal{U}})}$, $\|A\|_{C^{2}({\mathcal{U}})}$ and $C_{0}$ (the value of $C$ may change even in the same line).
Differentiating (21) with respect to $x_{1}$, we obtain
$w_{111}w_{22}=\phi_{1}-w_{11}w_{221}+2w_{12}w_{112}=\phi_{1}-w_{11}w_{22}\frac{w_{221}}{w_{22}}=\phi_{1}-\phi\frac{w_{221}}{w_{22}}.$
Whence, using again $w_{22}=\lambda_{2}\geq C_{0}$,
$|w_{111}|\leq C(\phi+|\nabla\phi|).$
A similar inequality can be deduced for $w_{112}$. Therefore
(23) $|\nabla w_{11}|\leq C(\phi+|\nabla\phi|).$
Our aim is to compute
$\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}(z)\phi_{\alpha\beta}(z).$
We start observing that, as a consequence of the Codazzi formula, the quantity
(24) $w_{ijk},\quad i,j,k=1,\dots,n,$
is independent of the order of $i,j,k$. We have, for $\alpha,\beta=1,2$,
$\phi_{\alpha\beta}$ $=$ $w_{11\alpha\beta}w_{22}+w_{11\alpha}w_{22\beta}+w_{11\beta}w_{22\alpha}+w_{11}w_{22\alpha\beta}-2(w_{12\alpha}w_{12\beta}+w_{12}w_{12\alpha\beta})$
Taking into account (22) and (23), we deduce the following expression for the second order derivatives of $\phi$:
(25) $\phi_{\alpha\beta}=w_{22}\left(w_{11\alpha\beta}-2\frac{1}{w_{22}}w_{12\alpha}w_{12\beta}\right)+O(\phi+|\nabla\phi|),\quad\alpha,\beta=1,2.$
Here and throughout the paper, by $O(\phi+|\nabla\phi|)$ we mean a quantity which is bounded in absolute value by a constant $C$ (as above) times $(\phi+|\nabla\phi|)$.
We also have the identities (see for instance [13]):
(26) $w_{ii\alpha\beta}=w_{\alpha\beta ii}+w_{ii}-w_{\alpha\beta},\quad\alpha,\beta,i=1,2,$
so that
$w_{11\alpha\beta}=w_{\alpha\beta 11}-w_{\alpha\beta}+w_{11}=w_{\alpha\beta 11}-w_{\alpha\beta}+O(\phi+|\nabla\phi|).$
Hence
$\phi_{\alpha\beta}=w_{22}w_{\alpha\beta 11}-2w_{12\alpha}w_{12\beta}-w_{22}w_{\alpha\beta}+O(\phi+|\nabla\phi|).$
Next we compute
$\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}\phi_{\alpha\beta}=w_{22}\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}w_{\alpha\beta 11}-2a_{22}w_{221}^{2}-w_{22}\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}w_{\alpha\beta}+O(\phi+|\nabla\phi|),$
or, equivalently,
(27) $\frac{1}{w_{22}}\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}\phi_{\alpha\beta}=-\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}w_{\alpha\beta 11}+2\frac{a_{22}w_{221}^{2}}{w_{22}}+a_{22}w_{22}+O(\phi+|\nabla\phi|).$
Here, we have used (22) to absorb the term $a_{11}w_{11}$ into $O(\phi+|\nabla\phi|)$.
Next, we differentiate twice equation (13) with respect to $x_{1}$, and we get:
$\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}w_{\alpha\beta 11}$ $=$ $f_{11}-2\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta 1}w_{\alpha\beta 1}-\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta 11}w_{\alpha\beta}$
$=$ $f_{11}-2a_{221}w_{221}-a_{2211}w_{22}+O(\phi+|\nabla\phi|),$
where $a_{\alpha\beta 1}$ and $a_{\alpha\beta 11}$ denote the first and the second partial derivatives of $a_{\alpha\beta}$ with respect to $x_{1}$, respectively, and we used (22) and (23). Replacing the last equality in (27) we obtain:
$-\frac{1}{w_{22}}\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}\phi_{\alpha\beta}$ $=$ $-f_{11}+2a_{221}w_{221}+2a_{2211}w_{22}+2\frac{a_{22}w_{221}^{2}}{w_{22}}+a_{22}w_{22}+O(\phi+|\nabla\phi|)$
$=$ $f-f_{11}+2\frac{w_{221}}{w_{22}}(a_{22}w_{22})_{1}+a_{2211}w_{22}+O(\phi+|\nabla\phi|).$
Using again the equation (13), we obtain
$w_{22}=\frac{f}{a_{22}}-\frac{a_{11}}{a_{22}}w_{11}=\frac{f}{a_{22}}+O(\phi+|\nabla\phi|).$
Here we used (22) and the fact that, due to the positivity of the tensor $A$, $a_{22}\geq c$ for some constant $c>0$ depending on $A$ only. Moreover, differentiating once (13) with respect to $x_{1}$ and taking into account (22) and (23), we get
$a_{221}w_{22}+a_{22}w_{221}=f_{1}+O(\phi+|\nabla\phi|).$
Therefore
$\frac{w_{221}}{w_{22}}=\frac{f_{1}}{f}-\frac{a_{221}}{a_{22}}+O(\phi+|\nabla\phi|).$
For simplicity, let $a=a_{22}$, $a_{1}=a_{221}$ and $a_{11}=a_{2211}$. From the previous steps we deduce
$-\frac{1}{w_{22}}\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}\phi_{\alpha\beta}$ $=$ $f-f_{11}+2\frac{f_{1}^{2}}{f}-2\frac{a_{1}f_{1}}{a}+\frac{fa_{11}}{a}+O(\phi+|\nabla\phi|)$
$=$ $\frac{f^{2}}{a}\left[\left(\frac{a}{f}\right)_{11}+\frac{a}{f}\right]+O(\phi+|\nabla\phi|).$
By (16), the first term in the right hand side of the previous relation is non-negative:
(28) $\left(\frac{a}{f}\right)_{11}+\frac{a}{f}\geq 0.$
Hence, we have the following differential inequality in $\mathcal{U}$:
$\sum_{\alpha,\beta=1}^{2}a_{\alpha\beta}\phi_{\alpha\beta}\leq C(\phi+|\nabla\phi|),$
for every $z\in{\mathcal{U}}$. As $\phi\geq 0$ and $\phi(x_{0})=0$, it follows from the strong minimum principle that $\phi\equiv 0$ in $D$, as $D$ is connected. This is equivalent to say that the rank of $W$ is constantly $1$ in $D$.
The proof of the final part of the statement, relative to the case $D=\mathbb{S}^{2}$, is provided in Section 3, in the more general setting of arbitrary dimension $n$.
2.2. Proof of Theorems 1.2, 1.4 and 1.5
We will need the following two remarks.
Remark 2.1.
Let
$M\colon\mathbb{R}^{n+1}\to\mbox{\rm Sym}(n),$
and let $M_{ij}$ denote the $ij$-th entry of $M$, for $i,j=1,\dots,n$. If $M$ is convex, then in particular $M_{ii}$ is a convex function in $\mathbb{R}^{n+1}$, for every $i=1,\dots,n+1$.
Remark 2.2.
Let $g\colon{{\mathbb{S}}^{n}}\to\mathbb{R}$ and let $G\colon\mathbb{R}^{n+1}\to\mathbb{R}$ be the $1$-homogeneous extension of $g$. Assume that $g\in C^{2}({{\mathbb{S}}^{n}})$, and let $g_{ij}$, $i,j=1,\dots,n$, denote the second covariant derivatives of $g$, with respect to a local orthonormal frame $\{\bar{e}_{1},\dots,\bar{e}_{n}\}$ on ${\mathbb{S}}^{n}$}, defined on some open neighborhood of a point $x_{0}\in{{\mathbb{S}}^{n}}$. We now choose an orthonormal frame $\{e_{1},\dots,e_{n},e_{n+1}\}$ in $\mathbb{R}^{n+1}$, such that $\{e_{1},\dots,e_{n}\}$ coincides with $\{\bar{e}_{1},\dots,\bar{e}_{n}\}$ at $x_{0}$. Then
$g_{ij}(x_{0})+g(x_{0})\delta_{ij}=G_{ij}(x_{0})\quad\forall\,i,j=1,\dots,n,$
where $G_{ij}$, $i,j=1,\dots,n$, denote the Euclidean second derivatives of $G$. A proof of this fact can be found, for instance, in [11, Appendix].
Proof of Theorem 1.2.
We need to prove that condition (17) implies (16). Let us fix an orthonormal frame on ${\mathbb{S}}^{n}$}, and $q\in\{1,\dots,n\}$. By Remark 2.1, the convexity of $M$ implies that the $1$-homogeneous extension of $(a_{qq}/f)$ to $\mathbb{R}^{3}$ is convex. Then (16) follows from Remark 2.2. ∎
Proof of Theorem 1.5.
By (12), the tensor $A=(a_{ij})_{i,j=1,2}$ appearing in (13) is given by
$A=C(W_{0}),$
where $C(W_{0})$ is the co-factor matrix of $W_{0}$. Let
$W_{0}=\left(\begin{array}[]{cc}b_{11}&b_{12}\\ b_{21}&b_{22}\end{array}\right).$
Then, by (11) we see that:
$(a_{ij})_{i,j=1,2}=\left(\begin{array}[]{cc}b_{22}&-b_{12}\\ -b_{21}&b_{11}\end{array}\right)$
In particular, in this case the matrix
$\frac{1}{f}(a_{ij})_{i,j=1,2}$
has the same diagonal entries of the matrix
$\frac{1}{f}W_{0},$
up to a permutation. The convexity of the $1$-homogeneous extension of the latter matrix then implies that assumption (16) is verified, by the same argument used in the proof of Theorem 1.2. ∎
Proof of Theorem 1.4.
The proof is based on an inspection of the proofs of Theorem 1.1 in the case $n=2$, and of the proof of Theorem 1.5. In the final part of the proof of Theorem 1.1 for $n=2$, we have seen that the argument of the proof works under the condition:
$\left(\frac{a_{22}}{f}\right)_{11}+\frac{a_{22}}{f}\geq 0.$
Here, a local orthonormal coordinate system $\{e_{1},e_{2}\}$ has been fixed on $\mathbb{S}^{2}$, and the subscript $1$ denotes the covariant differentiation along $e_{1}$. On the other hand, as we have seen in the proof of Theorem 1.5,
$a_{22}=w_{11}=h_{11}+h,$
where $h$ is the support function of the given convex body $\Omega_{0}$. Hence the condition becomes
$\left(\frac{h_{11}+h}{f}\right)_{11}+\frac{h_{11}+h}{f}\geq 0.$
Note that the direction $e_{1}$ may be completely arbitrary. We conclude that in Theorem 1.1, (16) can be replaced by condition:
(29) $\left(\frac{h_{\alpha\alpha}(x)+h(x)}{f(x)}\right)_{\alpha\alpha}+\frac{h_{\alpha\alpha}(x)+h(x)}{f(x)}\geq 0,\quad\forall\,x\in\mathbb{S}^{2},\,\forall\,\alpha,$
where $\alpha$ is a tangential direction to $\mathbb{S}^{2}$ at $x$. ∎
3. The $n$-dimensional case
This section is devoted to the proof of Theorem 1.1, when $n\geq 3$.
3.1. Some preliminary facts
We assume throughout that $u$ is a solution to (13) in $D$, and that the assumptions of Theorem 1.1 concerning the tensor $(a_{\alpha\beta})_{\alpha,\beta=1,\dots,n}$ and the function $f$, are verified. The tensor $W=W(x)$ is defined in the statement of the theorem; as usual $w_{ij}$, $i,j=1,\dots,n$, denote the entries of $W$. By assumption (14) and standard regularity theory of elliptic equations, we have $u\in C^{4,\gamma}({{\mathbb{S}}^{n}})$, for some $\gamma\in(0,1)$.
Given a square symmetric matrix $M\in\mbox{Sym}(n)$ and $j\in\{0,1,\dots,n\}$, let $\sigma_{j}(M)$ denote the $j$th elementary symmetric function of the eigenvalues of $M$. Note that, if $M$ is positive semi-definite, then the rank of $M$ is $l\in\{0,\dots,n-1\}$ if and only if
$\sigma_{l}(M)>0\quad\mbox{and}\quad\sigma_{j}(M)=0\quad\forall\,j=l+1,\dots,n.$
If
$\sigma_{n}(W(x))>0\quad\forall\,x\in D,$
then $W$ has constant (and maximal) rank $n$ in $D$. So, in the perspective to prove Theorem 1.1, we will assume that there exists an integer $l$ between 0 and (n1)(n-1), with the following property:
$\sigma_{l}(W(x))>0,\quad\forall\,x\in D,$
and there exists $x_{0}\in D$, such that
$\sigma_{l+1}(W(x_{0}))=0.$
Remark 3.1.
As already observed in the two-dimensional case, $l$ is at least $1$ by equation (13) and the strict positivity of $f$.
We now introduce the test function $\phi$, which is one of the main tools of this proof. We remark that the choice of $\phi$ for $n\geq 3$ is different from the one in the case $n=2$. Define
(30) $\phi(x)=\sigma_{l+1}(W(x))+\frac{\sigma_{l+2}(W(x))}{\sigma_{l+1}(W(x))},$
for every $x$ such that $\sigma_{l+1}(W(x))>0$.
Remark 3.2.
This test function was introduced in [4], where the authors observed that, by the Newton-McLaurin inequalities, $\phi$ can be continuously extended to the case $\sigma_{l+1}(W)=0$, setting
$\frac{\sigma_{l+2}(W)}{\sigma_{l+1}(W)}=0\quad\mbox{if $\sigma_{l+1}(W)=0$.}$
Therefore we complete the definition of $\phi$ setting
(31) $\phi(x)=0,\quad\mbox{$\forall\,x$ s.t. $\sigma_{l+1}(W(x))=0$.}$
The general strategy of the proof will be to show that, under the assumptions of the theorem, $\phi$ verifies a differential inequality, to which the strong minimum principle can be applied. This will lead to the conclusion
$\phi\equiv 0\quad\mbox{in $D$,}$
that is, $W$ has constant rank $l$ in $D$.
Let $D^{\prime}$ be an open and connected set compactly contained in $D$, and let $C_{0}>0$ be such that
(32) $\sigma_{l}(W)\geq C_{0}\quad\mbox{in $D^{\prime}$.}$
For any $z\in D^{\prime}$, let $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ be the eigenvalues of the matrix $W$ at $z$. As $u\in C^{3}(D)$, there exists a constant $C>0$, depending on $\|u\|_{C^{3}}$, $\|\phi\|_{C^{2}}$, $C_{0}$ and $n$, such that
$\lambda_{n}\geq\lambda_{n-1}\geq\dots\geq\lambda_{n-l+1}\geq C,\quad\mbox{in $D^{\prime}$}.$
We set
$B=\{1,\dots,n-l\},\quad G=\{n-l+1,\dots,n\}.$
We now recall a notation which is common in the context of constant rank type results. Given two functions $g$ and $h$ defined in $D^{\prime}$, and $y\in D^{\prime}$, we write
$g(y)\lesssim h(y)$
if there exist positive constants $c_{1}$ and $c_{2}$, such that
$g(y)-h(y)\leq c_{1}|\nabla\phi(y)|+c_{2}\phi(y).$
We also write
$g(y)\sim h(y)$
if $g(y)\lesssim h(y)$ and $g(y)\lesssim h(y)$. Moreover, we write
$g\lesssim h\qquad\mbox{in $D^{\prime}$}$
if this inequality holds with constants $c_{1}$ and $c_{2}$ independent of $y$ and depending only on $\|u\|_{C^{3}}(D^{\prime})$, $\|\phi\|_{C^{2}}(D^{\prime})$, $C_{0}$ and $n$.
As in the proof of Theorem 1.1 in the case $n=2$, we also write
$g=O(\psi),$
for some positive function $\psi$ defined in $D^{\prime}$, if
$|g|\leq C\psi\quad\mbox{in $D^{\prime}$,}$
where, as above, the constant $C$ depends only on $\|u\|_{C^{3}}$, $\|\phi\|_{C^{2}}$, $C_{0}$ and $n$.
In view of the Codazzi property of $W$ and arguments presented in [4], $\phi$ has the properties listed below.
  1. I.
    Once extended as indicated in (31), $\phi\in C^{1,1}$ (see Corollary 2.2 in [4]).
  2. II.
    $\forall\alpha,\beta=1,\dots,n$,
    (33) $\phi_{\alpha\beta}$ $=$ $O\left(\phi+\sum_{i,j\in B}|\nabla w_{ij}|\right)-\left[\frac{\sum_{i\in B}V_{i\alpha}V_{i\beta}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}w_{ij\alpha}w_{ji\beta}}{\sigma_{1}(B)}\right]+$
    $+\sum_{i\in B}\left[\sigma_{l}(G)+\frac{\sigma_{1}^{2}(B|i)+\sigma_{2}(B|i)}{\sigma_{1}^{2}(B)}\right]\left[w_{ii\alpha\beta}-2\sum_{j\in G}\frac{w_{ji\alpha}w_{ij\beta}}{w_{jj}}\right],$
    where
    $V_{i\alpha}=\sigma_{1}(B)w_{ii\alpha}-w_{ii}\sum_{j\in B}w_{jj\alpha},$
    and $\sigma_{1}(B)$, $\sigma_{k}(B|i)$ denote respectively the 11st, $k$th elementary symmetric function of the eigenvalues with indices in $B$, $B\setminus\{i\}$. Relation (33) follows from Lemma 2.4 and Proposition 2.1 in [4] (see also the argument used to deduce (3.19) there).
  3. III.
    Let m=|B|=nlm=|B|=n-l. By Lemma 3.3 in [4], there is a constant $C$ depending only on $n$, $\|W\|_{C^{0}}$ and $C_{0}$ in (32), such that for each $\alpha=1,\dots,n$, for any $K>0$, $\delta>0$,
    (34) $\sum_{i,j=1}^{m}|w_{ij\alpha}|\leq C\left(1+\frac{2}{\delta}+K\right)(\phi+|\nabla\phi|)+\frac{\delta}{2K}\frac{\sum_{i\neq j}^{m}|w_{ij\alpha}|^{2}}{\sigma_{1}(B)}+\frac{C}{K}\frac{\sum_{i=1}^{m}V_{i\alpha}^{2}}{\sigma_{1}^{3}(B)}.$
Remark $3.$
By (33), (34) and the uniform ellipticity of $(a_{\alpha\beta})$, in the rest of the proof, we may write $w_{\alpha\beta i}\sim 0$ if $i\in B$ and at least one of $\alpha,\beta$ is in $B$.
3.2. The differential inequality for the function $\phi$
In the sequel we will use the same notation adopted in the proof for the case $n=2$, for the covariant derivatives of the coefficients $a_{\alpha\beta}$. In particular,
$a_{\alpha\beta i},\quad a_{\alpha\beta ij}$
denote the first and second covariant derivatives of $a_{\alpha\beta}$, along the directions $e_{i}$, and along the directions $e_{i}$ and $e_{j}$ respectively, of a local orthonormal coordinate system on ${\mathbb{S}}^{n}$}.
Proposition 3.4.
Let $D^{\prime}$ be an open subset of $\mathbb{S}^{n}$, and let $u\in C^{3}(\Omega)$ be a solution of equation (13) in $D^{\prime}$, such that the matrix $W$ defined in (1.1) is positive semi-definite in $D^{\prime}$. Then $\phi$ verifies the following relation in $D^{\prime}$
$\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}\phi_{\alpha\beta}$ $\leq$ $\sum_{i\in B}\left[\sigma_{l}(G)+\frac{\sigma_{1}^{2}(B|i)+\sigma_{2}(B|i)}{\sigma_{1}^{2}(B)}\right]\times$
$\times\left(-f+f_{ii}-\sum_{\alpha\in G}a_{\alpha\alpha ii}w_{\alpha\alpha}-2\sum_{\alpha,\beta\in G}\left[a_{\alpha\beta i}w_{\alpha\beta i}+\sum_{j\in G}\frac{a_{\alpha\beta}w_{ij\alpha}w_{ij\beta}}{w_{jj}}\right]\right)$
$+O(\phi+|\nabla\phi|).$
Proof.
For an arbitrary $z\in D^{\prime}$, we choose a local orthonormal frame $\{e_{1},\dots,e_{n}\}$ defined in a neighborhood of $z$, and we assume that $W(z)$ is diagonal, with the convention on sets $G$ and $B$ of “good” and “bad” subsets of indices of the eigenvalues of $W$ described before. We have
$w_{ii}\sim 0\quad\forall\,i\in B,\quad w_{ij}\sim 0\quad\mbox{for $i\neq j$.}$
As a consequence of (33) and (26), we obtain
$\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}\phi_{\alpha\beta}$ $\leq$ $\sum_{i\in B}\left[\sigma_{l}(G)+\frac{\sigma_{1}^{2}(B|i)+\sigma_{2}(B|i)}{\sigma_{1}^{2}(B)}\right]\left[\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta ii}\right$.
$\left.-\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta}-2\sum_{\alpha,\beta=1}^{n}\sum_{j\in G}a_{\alpha\beta}\frac{w_{ij\alpha}w_{ij\beta}}{w_{jj}}\right]$
$-\sum_{\alpha,\beta}\left[\frac{\sum_{i\in B}a_{\alpha\beta}V_{i\alpha}V_{i\beta}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}a_{\alpha\beta}w_{ij\alpha}w_{ji\beta}}{\sigma_{1}(B)}\right]$
$+O\left(\phi+\sum_{i,j\in B}|\nabla w_{ij}|\right).$
By the uniform ellipticity, there exists $\delta_{0}>0$ such that
$(a_{\alpha\beta})\geq\delta_{0}I_{n\times n},$
where $I_{n\times n}$ is the identity matrix of order $n$. Therefore
$\sum_{\alpha,\beta}\left[\frac{\sum_{i\in B}a_{\alpha\beta}V_{i\alpha}V_{i\beta}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}a_{\alpha\beta}w_{ij\alpha}w_{ji\beta}}{\sigma_{1}(B)}\right]\geq\delta_{0}\left[\frac{\sum_{i\in B}V^{2}_{i\alpha}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}|w_{ij\alpha}|^{2}}{\sigma_{1}(B)}\right]$
Therefore, by (3.2),
$\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}\phi_{\alpha\beta}$ $\leq$ $\sum_{i\in B}\left[\sigma_{l}(G)+\frac{\sigma_{1}^{2}(B|i)+\sigma_{2}(B|i)}{\sigma_{1}^{2}(B)}\right]\left[\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta ii}\right$.
$\left.-\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta}-2\sum_{\alpha,\beta=1}^{n}\sum_{j\in G}a_{\alpha\beta}\frac{w_{ij\alpha}w_{ij\beta}}{w_{jj}}\right]$
$-\delta_{0}\left[\frac{\sum_{i\in B}V^{2}_{i\alpha}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}|w_{ij\alpha}|^{2}}{\sigma_{1}(B)}\right]+O\left(\phi+\sum_{i,j\in B}|\nabla w_{ij}|\right).$
Next, we use (34) to deduce that
$\sum_{i,j\in B}|\nabla w_{ij}|$
can be controlled by
$\delta_{0}\left[\frac{\sum_{i\in B}V^{2}_{i\alpha}}{\sigma^{3}_{1}(B)}+\frac{\sum_{i,j\in B,i\neq j}|w_{ij\alpha}|^{2}}{\sigma_{1}(B)}\right]+O(\phi+|\nabla\phi|).$
Hence
$\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}\phi_{\alpha\beta}$ $\leq$ $\sum_{i\in B}\left[\sigma_{l}(G)+\frac{\sigma_{1}^{2}(B|i)+\sigma_{2}(B|i)}{\sigma_{1}^{2}(B)}\right]\left[\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta ii}\right$.
$\left.-\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}w_{\alpha\beta}-2\sum_{\alpha,\beta=1}^{n}\sum_{j\in G}a_{\alpha\beta}\frac{w_{ij\alpha}w_{ij\beta}}{w_{jj}}\right]$
$+O(\phi+|\nabla\phi|).$
Next, for $i\in B$, we differentiate twice (13) with respect to $e_{i}$:
$\sum_{\alpha,\beta=1}^{n}\left[a_{\alpha\beta ii}w_{\alpha\beta}+2a_{\alpha\beta i}w_{\alpha\beta i}+a_{\alpha\beta}w_{\alpha\beta ii}\right]=f_{ii};$
whence
$\sum_{\alpha,\beta=1}^{n}\sum_{i\in B}a_{\alpha\beta}w_{\alpha\beta ii}$ $=$ $\sum_{i\in B}f_{ii}-\sum_{\alpha,\beta=1}^{n}\sum_{i\in B}a_{\alpha\beta ii}w_{\alpha\beta}-2\sum_{\alpha,\beta=1}^{n}\sum_{i\in B}a_{\alpha\beta i}w_{\alpha\beta i}$
$\sim$ $\sum_{i\in B}f_{ii}-\sum_{\alpha\in G}^{n}\sum_{i\in B}a_{\alpha\alpha ii}w_{\alpha\alpha}-2\sum_{\alpha,\beta=1}^{n}\sum_{i\in B}a_{\alpha\beta i}w_{\alpha\beta i}.$
Replacing this relation into (3.2) we obtain (3.4). ∎
We will need a further auxiliary result. Let ${\mathcal{M}}_{l\times l}$ be the space of $l\times l$ matrices, where $l\in\mathbb{N}$. We consider the standard Euclidean inner product for matrices defined by
$\langle M_{1},M_{2}\rangle=\mbox{tr}(M_{1}M_{2}^{T}),$
where $M_{1},M_{2}\in{\mathcal{M}}_{l\times l}$, and the apex $T$ denotes the transpose.
Lemma 3.5.
Let $l\geq 2$ be an integer number and let $R,S,T\in{\mathcal{M}}_{l\times l}$. Assume in particular that $R$ is symmetric and positive definite. Let
${\mathcal{D}}=\{W\in\mbox{\rm Sym}(l)\colon\langle R,W\rangle=1,W\geq 0\}.$
Consider the function $\rho\colon{\mathcal{D}}\to\mathbb{R}$ defined by
$\rho(W)=\langle S,W\rangle+\langle T,W\rangle^{2}.$
Then there exists $W_{\rm min}\in\mathcal{D}$, with $rank(W_{\rm min})=1$, such that
$\rho(W_{\rm min})=\min_{\mathcal{D}}\rho.$
Proof.
Note that $\mathcal{D}$ is convex and compact (the second claim follows being $R$ positive definite), and that $\langle R,R\rangle>0$. Let
$S^{\prime}=S-\frac{{\langle S,R\rangle}}{\langle R,R\rangle}R,\quad T^{\prime}=T-\frac{{\langle T,R\rangle}}{\langle R,R\rangle}R.$
Clearly
$\langle S^{\prime},R\rangle=\langle T^{\prime},R\rangle=0.$
Now, using that $\langle R,W\rangle=1$, $\rho(W)$ rewrites for $W\in{\mathcal{D}}$ as
$\rho(W)$ $=$ $\langle S^{\prime},W\rangle+\frac{\langle S,R\rangle}{\langle R,R\rangle}+\left[\langle T^{\prime},W\rangle+\frac{\langle T,R\rangle}{\langle R,R\rangle}\right]^{2}$
$=$ $\langle S^{\prime},W\rangle+a+\left[\langle T^{\prime},W\rangle+b\right]^{2},$
where $a$ and $b$ are constants. Hence, it suffices to prove the lemma for the function $\bar{\rho}$ defined by
$\bar{\rho}(W)=\langle S^{\prime},W\rangle+\left[\langle T^{\prime},W\rangle+b\right]^{2},\quad W\in{\mathcal{D}}.$
As $\mathcal{D}$ is compact, there exists $W_{0}\in{\mathcal{D}}$, such that
$\bar{\rho}(W_{0})=\min_{W\in\mathcal{D}}\bar{\rho}(W).$
Let $k$ be the rank of $W_{0}$, so $1\leq k\leq l$. If $k=1$, then the conclusion of the lemma is true. Hence consider the case $k\geq 2$. After a rotation, we may assume that the entries $w_{ij}$ of $W_{0}$ verify $w_{ij}=0$ for every $i>k$. We may also embed $\mbox{\rm Sym}(k)$ in $\mbox{\rm Sym}(l)$, with the convention that
$\mbox{\rm Sym}(k)=\{W=(w_{ij})_{i,j=1,\dots,l}\in\mbox{\rm Sym}(l)|w_{ij}=0,\forall i>k\}.$
Define
${\mathcal{D}}_{k}=\{W\in\mbox{\rm Sym}(k)\colon\langle R,W\rangle=1,W\geq 0\}.$
Note that ${\mathcal{D}}_{k}\subset{\mathcal{D}}$ is convex and compact. Let $\mathcal{H}$ be the subspace of $\mbox{\rm Sym}(k)$ defined by
${\mathcal{H}}=\{Y\in\mbox{\rm Sym}(k)\colon\langle Y,R\rangle=0\}.$
In particular, $T^{\prime}\in{\mathcal{H}}$; moreover, $\mathcal{H}$} has dimension
$\frac{k(k+1)}{2}-1\geq 2,$
as $k\geq 2$. Hence there exists $\bar{T}\in{\mathcal{H}}$ such that
$\langle T^{\prime},\bar{T}\rangle=0,\quad\langle\bar{T},\bar{T}\rangle=1.$
As $\bar{T}\in{\mathcal{H}}$, the points of the form $W=W_{0}+x\bar{T}$, for $x\in\mathbb{R}$, verify $\langle W,R\rangle=1$. Let
$L=\{W\in\mbox{\rm Sym}(k)\colon W=W_{0}+x\bar{T}\}\quad\mbox{and}\quad\Sigma=L\cap{\mathcal{D}}.$
$L$ is a straight line in $\mbox{\rm Sym}(l)$, lying on the same hyperplane as ${\mathcal{D}}_{k}$, and by the convexity of ${\mathcal{D}}_{k}$, $\Sigma$ is a segment containing $W_{0}$ in its interior (because the rank of $W_{0}$ is $k$), that is
$\Sigma=\{W_{0}+x\bar{T}\colon x\in[-c,d]\}$
for some $c,d>0$. In particular $W_{0}-c\bar{T}$ and $W_{0}+d\bar{T}$ belong to the relative boundary of ${\mathcal{D}}_{k}$ (meaning that they are points of ${\mathcal{D}}_{k}$ not belonging to the relative interior of ${\mathcal{D}}_{k}$). Consider the function $\gamma\colon[-c,d]\to\mathbb{R}$ defined by
$\gamma(x)$ $=$ $\bar{\rho}(W_{0}+x\bar{T})=\langle W_{0}+x\bar{T},S^{\prime}\rangle+\left[\langle W_{0}+x\bar{T},T^{\prime}\rangle+b\right]^{2}$
$=$ $\langle W_{0}+x\bar{T},S^{\prime}\rangle+\left[\langle W_{0},T^{\prime}\rangle+b\right]^{2}.$
Thus, $\gamma$ is a linear function with a minimum point at $x=0$. Therefore $\gamma$ is constant, which shows that the minimum of $\bar{\rho}$ is attained (also) on the relative boundary of ${\mathcal{D}}_{k}$. In other words, there exists $\tilde{W}_{0}\in{\mathcal{D}}_{k-1}$ such that
$\bar{\rho}(\tilde{W}_{0})=\min_{W\in{\mathcal{D}}}\bar{\rho}(W).$
Iterating the previous argument, we deduce that the minimum is attained on matrices with rank one. ∎
Proof of Theorem 1.1..
As $f>0$, we may assume that $f\equiv 1$ on $D$, up to passing from the tensor $A$ to
$\frac{A}{f}.$
Let $l\in\{1,\dots,n-1\}$ be the minimal rank of the matrix $W$ on $D$, attained at some point $x_{0}\in D$. Now fix any point $z$ in a sufficiently small neighborhood $D^{\prime}$ of $x_{0}$ and call $\ell$ the rank of $W$ at $z$. We have $\ell\geq l$. We will consider $W=W(z)$ as an $\ell\times\ell$ matrix, with the convention that $w_{ij}=0$, for every $i>\ell$. We set
$(w^{jk})_{j,k=1,\dots,\ell}=W^{-1},$
if no ambiguity may occur. The next calculation will be performed at the point $z$.
We will prove that, in $D^{\prime}$, for every $i\in B$,
(39) $\sum_{\alpha,\beta=1}^{\ell}a_{\alpha\beta ii}w_{\alpha\beta}+2\sum_{\alpha,\beta=1}^{\ell}\left(a_{\alpha\beta i}w_{\alpha\beta i}+\sum_{j,k=1}^{\ell}a_{\alpha\beta}w^{jk}w_{\alpha ki}w_{\beta ji}\right)\gtrsim-1.$
By (3.4), with $f\equiv 1$, the previous inequality implies that
(40) $\sum_{\alpha,\beta=1}^{n}a_{\alpha\beta}\phi_{\alpha\beta}\lesssim 0.$
As a consequence of (40) and the strong minimum principle, $\phi\equiv 0$ in $D$, and the first conclusion of Theorem 1.1 will follow.
Let us fix $i\in B=\{\ell+1,\dots,n\}$, and set
$\Gamma_{i}:=\sum_{\alpha,\beta=1}^{\ell}a_{\alpha\beta ii}w_{\alpha\beta}+2\sum_{\alpha,\beta=1}^{\ell}\left(a_{\alpha\beta i}w_{\alpha\beta i}+\sum_{j,k=1}^{\ell}a_{\alpha\beta}w^{jk}w_{\alpha ki}w_{\beta ji}\right).$
In the rest of the proof, for simplicity we use == for $\sim$. The following relations hold:
(41) $\sum_{\alpha,\beta=1}^{\ell}a_{\alpha\beta}w_{\alpha\beta}=1,\quad\sum_{\alpha,\beta=1}^{\ell}a_{\alpha\beta i}w_{\alpha\beta}=-\sum_{\alpha,\beta=1}^{\ell}a_{\alpha\beta}w_{\alpha\beta i}.$
Here we used the fact that, by Remark 3.3, $w_{\alpha\beta i}\sim 0$ if at least one of $\alpha,\beta$ is in $B$. We set:
$A=(a_{\alpha\beta})_{\alpha,\beta=1}^{\ell},\;W=(w_{\alpha\beta})_{\alpha,\beta=1}^{\ell},\;A^{\prime}=(a_{\alpha\beta i})_{\alpha,\beta=1}^{\ell},\;A^{\prime\prime}=(a_{\alpha\beta ii})_{\alpha,\beta=1}^{\ell},\;W^{\prime}=(w_{\alpha\beta i})_{\alpha,\beta=1}^{\ell};$
note that $W^{\prime}$ is symmetric. Then $\Gamma_{i}$ can be written in the form
(42) $\Gamma_{i}(W,W^{\prime})=\langle A^{\prime\prime},W\rangle+2\left[\langle A^{\prime},W^{\prime}\rangle+\langle W^{\prime},\Theta W^{\prime}\rangle\right],$
where
$\Theta W^{\prime}=W^{-1}W^{\prime}A,$
and the equalities in (41) become
(43) $\langle A,W\rangle=1,\quad\langle A^{\prime},W\rangle=-\langle A,W^{\prime}\rangle.$
In the above constraints, == should be intended as $\sim$, in the sense that the right hand side of the first one should be $1+a(z)$, with $a(z)\sim 0$, and in the right hand side of the second one we should add a term $b(z)$, with $b(z)\sim 0$. However, notice that this will not affect the conclusion, which is that we can calculate the minimum of a certain functional $\tilde{\Gamma}_{i}$ over a certain subset $\mathcal{C}$ of $\mbox{\rm Sym}(\ell)\times\mbox{\rm Sym}(\ell)$ at a matrix of rank 1, so that we get $\Gamma_{i}\gtrsim-1$, that is, (39). The set $\mathcal{C}$ should indeed depend on $z$, via $a(z)$ and $b(z)$, but the conclusion does not change, for every fixed $z$, as desired.
Now, with a slight abuse of notation, we consider two fixed arbitrary matrices $A$ and $A^{\prime}$ in $\mathcal{M}$$_{\ell\times\ell}$, such that $A$ is symmetric and positive definite, and we define
(44) ${\mathcal{C}}=\{(W,Y)\in\mbox{\rm Sym}(\ell)\times\mbox{\rm Sym}(\ell)\colon W\geq 0,\ \langle A,W\rangle=1,\ \langle A^{\prime},W\rangle=-\langle A,Y\rangle\}.$
Note that if $(W,Y)\in{\mathcal{C}}$ and $W$ is positive definite, then $(W,Y)$ belongs to the relative interior of $\mathcal{C}$. We consider now the function $\hat{\Gamma}_{i}\colon{\mathcal{C}}\to\mathbb{R}$, defined by:
(45) $\hat{\Gamma}_{i}(W,Y)=\langle A^{\prime\prime},W\rangle+2\left[\langle A^{\prime},Y\rangle+\langle Y,\Theta Y\rangle\right],$
where $\Theta\colon\mbox{\rm Sym}(\ell)\to\mbox{\rm Sym}(\ell)$ is the linear map defined by
$\Theta P=W^{-1}PA,\quad\forall\,P\in\mbox{\rm Sym}(\ell).$
The scope of this part of the proof is to minimize $\hat{\Gamma}_{i}$ over $\mathcal{C}$. If we identify $\mathcal{M}$$_{\ell\times\ell}$ with $\mathbb{R}^{\ell^{2}}$, $\langle P,Q\rangle$ is the canonical inner product in $\mathbb{R}^{\ell^{2}}$. Adopting this point of view, $\Theta$ can be seen as a positive definite matrix, to which we associate a linear operator from $\mathbb{R}^{\ell^{2}}$ to $\mathbb{R}^{\ell^{2}}$, and
(46) $\Theta^{-1}Q=WQA^{-1},\quad\forall\ Q\in\mathbb{R}^{\ell^{2}}.$
Let $(W_{m},Y_{m})$, $m\in\mathbb{N}$, be a minimizing sequence for $\hat{\Gamma}_{i}$ on $\mathcal{C}$, such that $W_{m}>0$ for every $m$. As $\Theta$ depends on $W$, we denote by $\Theta_{m}$ the operator corresponding to $W_{m}$.
It follows from the expression of $\hat{\Gamma}_{i}$ that, for each fixed $m$, the function
$X\rightarrow\hat{\Gamma}_{i}(W_{m},X),$
defined on the set $\{X\in{\mathcal{M}}_{\ell\times\ell}\colon(W_{m},X)\in\mathcal{C}\}$, has a unique minimum point $X_{m}$. By the Lagrange multipliers theorem, there exists $\lambda\in\mathbb{R}$ such that
$A^{\prime}+2\Theta_{m}X_{m}=\lambda A\quad\Rightarrow\quad 2\Theta_{m}X_{m}=\lambda A-A^{\prime}.$
Therefore, using the definition of $\Theta^{-1}$,
$X_{m}$ $=$ $\frac{1}{2}\Theta_{m}^{-1}(\lambda A-A^{\prime})$
$=$ $\frac{1}{2}W_{m}(\lambda A-A^{\prime})A^{-1}$
$=$ $\frac{1}{2}\lambda W_{m}-\frac{1}{2}W_{m}A^{\prime}A^{-1}.$
As, by the definition of $\mathcal{C}$,
$\langle A,X_{m}\rangle=-\langle W_{m},A^{\prime}\rangle\quad\mbox{and}\quad\langle A,W_{m}\rangle=1\quad\forall\,m,$
we obtain
$\langle W_{m},A^{\prime}\rangle=\frac{1}{2}\lambda-\frac{1}{2}\langle A,W_{m}A^{\prime}A^{-1}\rangle.$
On the other hand, since $A$ and $W_{m}$ are symmetric, we find
$\langle A,W_{m}A^{\prime}A^{-1}\rangle=\mbox{tr}(AW_{m}A^{\prime}A^{-1})=\mbox{tr}(W_{m}A^{\prime})=\langle W_{m},A^{\prime}\rangle.$
Therefore, $\lambda=-\langle W_{m},A^{\prime}\rangle$ so that
$X_{m}=-\frac{1}{2}\langle W_{m},A^{\prime}\rangle W_{m}-\frac{1}{2}W_{m}A^{\prime}A^{-1}=-\frac{1}{2}\langle W_{m},A^{\prime}\rangle W_{m}-\frac{1}{2}\Theta_{m}^{-1}A^{\prime}.$
By (45),
(47) $\hat{\Gamma}_{i}(W_{m},X_{m})=\langle A^{\prime\prime},W_{m}\rangle+2\bar{\Gamma}_{i}(W_{m},X_{m})$
where
$\bar{\Gamma}_{i}(W_{m},X_{m})=\langle A^{\prime},X_{m}\rangle+\langle X_{m},\Theta X_{m}\rangle.$
We compute now:
(48) $\bar{\Gamma}_{i}(W_{m},X_{m})$ $=$ $-\frac{1}{2}\langle A^{\prime},\langle W_{m},A^{\prime}\rangle W_{m}\rangle-\frac{1}{2}\langle A^{\prime},\Theta_{m}^{-1}A^{\prime}\rangle$
$+\frac{1}{4}\langle\langle W_{m},A^{\prime}\rangle W_{m}+\Theta_{m}^{-1}A^{\prime},\Theta_{m}\left(\langle W_{m},A^{\prime}\rangle W_{m}+\Theta_{m}^{-1}A^{\prime}\right)\rangle$
$=$ $-\frac{1}{2}\langle A^{\prime},W_{m}\rangle^{2}-\frac{1}{2}\langle A^{\prime},\Theta_{m}^{-1}A^{\prime}\rangle+\frac{1}{4}\langle\langle W_{m},A^{\prime}\rangle W_{m}+\Theta_{m}^{-1}A^{\prime},\langle W_{m},A^{\prime}\rangle\Theta_{m}W_{m}+A^{\prime}\rangle$
$=$ $-\frac{1}{2}\langle A^{\prime},W_{m}\rangle^{2}-\frac{1}{2}\langle A^{\prime},\Theta_{m}^{-1}A^{\prime}\rangle+\frac{1}{4}\langle\langle W_{m},A^{\prime}\rangle W_{m}+\Theta_{m}^{-1}A^{\prime},\langle W_{m},A^{\prime}\rangle A+A^{\prime}\rangle$
$=$ $-\frac{1}{2}\langle A^{\prime},W_{m}\rangle^{2}-\frac{1}{2}\langle A^{\prime},\Theta_{m}^{-1}A^{\prime}\rangle$
$+\frac{1}{4}(\langle A^{\prime},W_{m}\rangle^{2}\langle A,W_{m}\rangle+\langle A^{\prime},W_{m}\rangle^{2}+\langle\Theta_{m}^{-1}A^{\prime},A\rangle\langle A^{\prime},W_{m}\rangle+\langle\Theta_{m}^{-1}A^{\prime},A^{\prime}\rangle)$
$=$ $\frac{1}{4}\langle A^{\prime},W_{m}\rangle^{2}-\frac{1}{4}\langle A^{\prime},\Theta_{m}^{-1}A^{\prime}\rangle\,$
where in the last equality we have used that $\langle W_{m},A\rangle=1$ and that $\langle\Theta_{m}^{-1}A^{\prime},A\rangle=\langle W_{m},A^{\prime}\rangle$. Define
(49) $\tilde{\Gamma}_{i}(W)$ $=$ $\langle A^{\prime\prime},W\rangle+\frac{1}{2}\langle A^{\prime},W\rangle^{2}-\frac{1}{2}\langle A^{\prime},\Theta^{-1}A^{\prime}\rangle$
(50) $=$ $\langle A^{\prime\prime},W\rangle+\frac{1}{2}\langle A^{\prime},W\rangle^{2}-\frac{1}{2}\langle A^{\prime},WA^{\prime}A^{-1}\rangle.$
Note that, by (47) and (48) and the definition of $X_{m}$, we have $\tilde{\Gamma}_{i}(W_{m})=\hat{\Gamma}_{i}(W_{m},X_{m})\leq\hat{\Gamma}_{i}(W_{m},Y)$, for every $Y\in\{X\in{\mathcal{M}}_{\ell\times\ell}\colon(W_{m},X)\in\mathcal{C}\}$. We want to find the minimum of $\tilde{\Gamma}_{i}(W)$ under the constraint
$\langle A,W\rangle=1,\ W\geq 0.$
If $\ell\geq 2$, by Lemma 3.5, the minimum of $\tilde{\Gamma}_{i}$ is attained at some $W\geq 0$ with rank $1$. After a proper rotation, we may assume $W$ is diagonal with $w_{11}>0$ and $w_{jj}=0$ for every $j\geq 2$. From the constraint $\langle W,A\rangle=1$, we deduce
$w_{11}=\frac{1}{a_{11}}\,$,
whence
(51) $\min_{\langle A,W\rangle=1,\ W\geq 0}\tilde{\Gamma}_{i}(W)=\frac{a_{11}^{\prime\prime}}{a_{11}}+\frac{1}{2}\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\frac{1}{2}\sum_{i,j=1}^{\ell}\frac{a^{ij}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}},$
where
$(a^{ij})_{i,j=1,\dots,n}=A^{-1}.$
Therefore, by the assumption in (15) we conclude (39).
On the other hand, if $\ell=1$, then, obviously, as before the minimum of $\tilde{\Gamma}_{1}$ is attained at some $W\geq 0$ with rank $1$. After a proper rotation, we may assume $W$ is diagonal with $w_{11}>0$. Hence
(52) $\min_{\langle A,W\rangle=1,\ W\geq 0}\tilde{\Gamma}_{i}(W)=\frac{a_{11}^{\prime\prime}}{a_{11}}+\frac{1}{2}\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\frac{1}{2}\sum_{i,j=1}^{1}\frac{a^{ij}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}},$
and we conclude (39) as above by means of (15). In particular, note that in this case
$\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\sum_{i,j=1}^{1}\frac{a^{ij}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}}=0.$
This concludes the proof in the case of a general subset $D\subset{{\mathbb{S}}^{n}}$.
We now consider the case $D={{\mathbb{S}}^{n}}$ and we show that constant rank of $W$ on $\mathbb{S}^{n}$ implies that $W$ has full rank, i.e., $W>0$ on ${\mathbb{S}}^{n}$}. To this aim, we assume by contradiction that the (constant) rank is $l<n$; we know that $l\geq 1$. Thus,
$\sigma_{l}(W(x))=:q(x)>0,\quad\sigma_{l+1}(W(x))=0,\quad\forall x\in\mathbb{S}^{n}.$
Let $\delta_{0}>0$ be such that
$q(x)\geq\delta_{0},\quad\forall x\in\mathbb{S}^{n}.$
Let $\Omega$ be the convex body with support function $u$. We may assume $u\geq 0$ by moving the Steiner point of $\Omega$ to the origin. By Minkowski identities (see [21, Section 5.3.1]), there exists $C_{n,l}>0$, such that
$0$ $=$ $C_{n,l}\int_{\mathbb{S}^{n}}\sigma_{l+1}(W(x))dx$
$=$ $\int_{\mathbb{S}^{n}}u\sigma_{l}(W(x))dx$
$\geq$ $\int_{\mathbb{S}^{n}}uq(x)dx\geq\delta_{0}\int_{\mathbb{S}^{n}}udx.$
As $u\geq 0$, we deduce $u\equiv 0$, that is, a contradiction. ∎
Remark 3.6.
Let $A(x)=(a_{ij}(x))\in C^{2,\gamma}(\mathbb{S}^{n})$ be a positive $(2,0)$-tensor, for some $\alpha\in(0,1)$, and let $h\in C^{2,\alpha}({{\mathbb{S}}^{n}})$, $h>0$. Then the quantity
$\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\sum_{i,j=1}^{l}\frac{a^{i,j}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}}$
is invariant if we pass from $A$ to $hA$. In particular if the $1$-homogeneous extension of $A$ to $\mathbb{R}^{n}$ is convex, and if
$\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\sum_{i,j=1}^{l}\frac{a^{i,j}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}}\leq 0\quad\mbox{on ${{\mathbb{S}}^{n}}$,}$
then the assumptions of Theorem 1.1 are verified by $hA$, for every $h>0$.
Remark 3.7.
In the notations of the proof of Theorem 1.1, if $n=2$ and $l=1$, then
(53) $\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\sum_{i,j=1}^{1}\frac{a^{ij}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}}=0.$
In general,
(54) $\left(\frac{a^{\prime}_{11}}{a_{11}}\right)^{2}-\sum_{i,j=1}^{l}\frac{a^{ij}a^{\prime}_{1i}a^{\prime}_{1j}}{a_{11}}=0$
if $A^{\prime}=\lambda A$ for some $\lambda\in\mathbb{R}$.
Remark 3.8.
In general, for $l\geq 2$, if WA=AW,W>0WA=AW,\ W>0,
$\langle A^{\prime},W\rangle^{2}-\langle A^{\prime},WA^{\prime}A^{-1}\rangle\leq 0,$
with equality holds if and only if $A^{\prime}=\lambda A$ for some $\lambda\in\mathbb{R}$. To see this, we first observe that for any matrix $B$,
(55) $\langle B,B\rangle=\langle B,B^{T}\rangle$
where $B^{T}$ the transpose of $B$. Write
$A=A^{\frac{1}{2}}A^{\frac{1}{2}},\ W=W^{\frac{1}{2}}W^{\frac{1}{2}}.$
As $\langle A,W\rangle=1$, by (55)
$\langle A^{\frac{1}{2}}W^{\frac{1}{2}},A^{\frac{1}{2}}W^{\frac{1}{2}}\rangle=1.$
We also have
$\langle A^{\prime},WA^{\prime}A^{-1}\rangle=\langle W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}},W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}}\rangle,$
and
$\langle A^{\prime},W\rangle=\langle A^{\frac{1}{2}}W^{\frac{1}{2}},W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}}\rangle.$
By Cauchy-Schwarz inequality,
$\langle A^{\prime},W\rangle^{2}\leq\langle A^{\frac{1}{2}}W^{\frac{1}{2}},A^{\frac{1}{2}}W^{\frac{1}{2}}\rangle\langle W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}},W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}}\rangle=\langle W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}},W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}}\rangle,$
with equality holds if and only if
$W^{\frac{1}{2}}A^{\prime}A^{-\frac{1}{2}}=\lambda A^{\frac{1}{2}}W^{\frac{1}{2}}$
for some $\lambda\in\mathbb{R}$. As $WA=AW$, we get $A^{\prime}=\lambda A$. This is satisfied if $l=1$.

### v2

@@
 in $D$, then $W$ has constant rank in $D$. In particular, if $D=\mathbb{S}^{2}$ then $W$ has full rank in $\mathbb{S}^{2}$.
-Using the previous results, Theorem A and a continuity argument, we will deduce the following full rank theorem.
+Using a continuity argument, from Theorem 1.1 it can be deduced that the solution of equation (13) is in fact a geometric solution, that is, its $1$-homogeneous extension to $\mathbb{R}^{n}$ is convex.
+The application of the continuity argument requires additional assumptions, which are more restrictive in the case $n\geq 3$, due to the non-linearity of condition (15) with respect to the matrix $A$.
+For this reason we present the result in two distinct statements.
 Corollary 1.5.
-Under the same assumptions of Theorem 1.1 (resp. Theorem 1.2) on AA and $f$, with $D={{\mathbb{S}}^{n}}$, assume moreover that $f$ verifies (8). Let $u$ be a solution of equation (13). Then
+Let $n=2$. Under the same assumptions of Theorem 1.2 on AA and $f$, with $D=\mathbb{S}^{2}$, assume moreover that $f$ verifies (8). Let $u$ be a solution of equation (13). Then
+$W=(u_{ij}+u\delta_{ij})>0\quad\mbox{on $\mathbb{S}^{2}$.}$
+In particular, the $1$-homogeneous extension of $u$ to $\mathbb{R}^{n}$ is convex.
+Corollary 1.6.
+Let $n\geq 3$.
+Under the same assumptions of Theorem 1.1 on AA and $f$, with $D={{\mathbb{S}}^{n}}$, assume moreover that condition (15) is verified by the pair $f$ and $A_{t}$, for every $t\in[0,1]$, where
+$A_{t}:=(1-t)I_{n}+tA;$
+(here $I_{n}$ is the identity matrix of order $n$). Let $u$ be a solution of equation (13). Then
 $W=(u_{ij}+u\delta_{ij})>0\quad\mbox{on ${{\mathbb{S}}^{n}}$.}$
-As an application of Corollary 1.5, we get the following existence result for the mixed Minkowski problem.
+As an application of the previous corollaries we get sufficient conditions for the existence of a solution to the mixed Minkowski problem.
+In view of the difference between the cases $n=2$ and $n\geq 3$, we state them separately.
+In particular, the result in the case $n=2$ is stated in §1.4.
 Theorem 1.3.
+Let $n\geq 3$.
 Let $\Omega_{1},\dots,\Omega_{n}$ be convex bodies in $\mathbb{R}^{n+1}$, with boundary of class $C^{4,\gamma}$, for some $\gamma>0$, and positive Gauss curvature.
@@
 $A(x)=(a_{ij}(x))$
-be the mixed co-factor matrix of $W_{1}(x),\dots,W_{n-1}(x)$.
-Let $f\in C^{2,\gamma}(\mathbb{S}^{n})$, with $f>0$ on $\mathbb{S}^{n}$, and assume that $f$ verifies condition (7) and condition (8).
+be the mixed co-factor matrix of $W_{1}(x),\dots,W_{n-1}(x)$. Moreover, let
+$A_{t}:=(1-t)I_{n}+tA,\quad t\in[0,1]$
+Let $f\in C^{2,\gamma}(\mathbb{S}^{n})$, with $f>0$ on $\mathbb{S}^{n}$, and assume that $f$ verifies condition (7).
 Let $\mu$ be the Borel measure on ${\mathbb{S}}^{n}$} with density $f$ with respect to the (n1)(n-1)-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}.
-Suppose that, for any $1\leq l\leq n-1$, and any $q=1,\ldots,n$,
-(18) $\left(\delta_{ij}+\frac{\nabla_{i}\nabla_{j}(\frac{a_{qq}}{f})}{\frac{a_{qq}}{f}}+\frac{1}{2}\frac{\nabla_{i}a_{qq}\nabla_{j}a_{qq}}{a^{2}_{qq}}-\frac{1}{2}\sum_{\alpha,\beta=1}^{l}\frac{a^{\alpha\beta}\nabla_{i}a_{q\alpha}\nabla_{j}a_{q\beta}}{a_{qq}}\right)_{i,j=1,\dots,n}\geq 0.$
+Suppose that, for every $1\leq l\leq n-1$, for every $t\in[0,1]$ and for every $q=1,\ldots,n$, the condition
+(18) $\left(\delta_{ij}+\frac{\nabla_{i}\nabla_{j}(\frac{a_{qq}}{f})}{\frac{a_{qq}}{f}}+\frac{1}{2}\frac{\nabla_{i}a_{qq}\nabla_{j}a_{qq}}{a^{2}_{qq}}-\frac{1}{2}\sum_{\alpha,\beta=1}^{l}\frac{a^{\alpha\beta}\nabla_{i}a_{q\alpha}\nabla_{j}a_{q\beta}}{a_{qq}}\right)_{i,j=1,\dots,n}\geq 0$
+is verified by the matrix $A_{t}$, for every $t\in[0,1]$.
 Then there exists a convex body $\Omega$ in $\mathbb{R}^{n+1}$, of class $C^{2,+}$, such that
@@
 Moreover, $\Omega$ is uniquely determined up to translations.
-In the case $n=2$, condition (18) reads as (16).
 Remark 1.7.
@@
 $M(x)=|x|^{2}\left[f\left(\frac{x}{|x|}\right)\right]^{-1}D^{2}h(x).$
-The rest of the paper is divided in two distinct and almost independent parts. Section 2.1 is devoted to the three-dimensional case ($n=2$), while in the final section we consider the case $n\geq 3$.
+The rest of the paper is organized as follows. Section 2.1 is devoted to the proofs of all the results established in the three-dimensional case ($n=2$).
+In Section 3 we consider the case $n\geq 3$.
+The proof of Corollaries 1.5 and 1.6 is given in Section 4.
 Acknowledgments.
@@
 for some $\lambda\in\mathbb{R}$. As $WA=AW$, we get $A^{\prime}=\lambda A$. This is satisfied if $l=1$.
+4. Proof of Corollaries 1.5 and 1.6
+As stated in the introduction, we exploit a continuity argument. Let $f$ and $A$ be as in the statement of either Corollary 1.5 or Corollary 1.6. Let
+$A_{0}\equiv I_{n},\quad A_{1}=A,$
+where $I_{n}$ is the identity matrix. Moreover, for $t\in(0,1)$, let
+$A_{t}=(1-t)A_{0}+tA_{1}.$
+Note that, for every $t\in[0,1]$, $A_{t}$ and $f_{t}$ verify the assumptions of Theorem 1.1.
+For $n\geq 3$ this is directly assumed in Corollary 1.6.
+For $n=2$ this is due to the fact that condition (16) is linear with respect to $A$, and the condition is verified by the pair $f$, AA, and by the pair $f$, $I_{n}$ (here the assumption that condition (8) holds for $f$ is used – see Remark 1.7).
+For $t\in[0,1]$ let $u_{t}$ be the solution of equation (13), with $A=A_{t}$ and $f=f_{t}$, and let
+$W_{t}=((u_{t})_{ij}+u_{t}\delta_{ij})_{i,j=1,\dots,n}\geq 0,\quad t\in[0,1].$
+We consider the set
+${\mathcal{T}}=\{t\in[0,1]\colon W_{t}>0\,\mbox{on ${{\mathbb{S}}^{n}}$}\}.$
+By Theorem A, $0\in{\mathcal{T}}$, and in particular $\mathcal{T}\neq\emptyset$.
+Moreover, the condition $W_{t}>0$ implies that $\mathcal{T}$ is open.
+On the other hand, Theorem 1.1 ensures that $\mathcal{T}$ is closed.
+We deduce that ${\mathcal{T}}=[0,1]$, and this concludes the proof.